6. Find the area of the region bounded between the curves y = -x² + 4x + 7 and y = x² - 9 show all steps thx
Answer #1

| 100 The intersection points of curves y=-5c2+45177 and y=x²q -S2+4547=sRq 231²_48-16=0 x²-26-8=0 roots are, x=-2, 4 Asea bo